Gold has a work function of 4.9 eV. (a) Calculate the maximum kinetic energy, in joules, of the electrons emitted when gold is illuminated with ultraviolet radiation of frequency 1.7 x 10^15 Hz. (b) What is the energy expressed in eV?
(c) What is the stopping potential for these electrons? (e = - 1.6 x 10^19 - 6.6 x 10^34Js)
1
Expert's answer
2020-11-11T07:47:50-0500

"h\nu=\phi+E_{max}"

(a)

"E_{max}=h\nu-\phi=6.6\cdot10^{-34}\cdot1.7\cdot10^{15}-4.9\cdot1.6\cdot10^{-19}=3.38\cdot10^{-19}" J

(b)

"E_{max}=\frac{3.38\cdot10^{-19}}{1.6\cdot10^{-19}}=2.1" eV

(c)

"\varphi=\frac{E_{max}}{e}=\frac{3.38\cdot10^{-19}}{1.6\cdot10^{-19}}=2.1" V